# A Cognitive Evaluation Benchmark of Image Reasoning and Description for Large Vision-Language Models

### v3

@@ 3.3.1 Recognition
Table 2: Recognition score of LVLMs on Description task. For reference, the Recognition Score of Oracle is calculated based on the annotated [Description] in CogBench as an estimated upper bound.
@@
InstructBLIP-7B 0.40
Qwen-VL-Chat 0.42
LLaVA-v1.5-7B 0.40
LLaVA-v1.5-13B 0.41
mPLUG-Owl-2 0.37
ShareGPT4V-7B 0.47
ShareGPT4V-13B 0.49
GPT-4V 0.67
GPT-4o 0.72
Oracle 0.93
Table 2 shows the Recognition Score of models on Description task. GPT-4o achieves the best performance in terms of recognition, which means it can recognize and describe more entities. Besides, GPT-4o and GPT-4V are significantly better than other open-source LVLMs, indicating these open-source LVLMs still have some room for development before reaching the recognition capability of GPT-4. ShareGPT4V-7B and ShareGPT4V-13B achieves better performance than other open-source LVLMs. As it follows the design of LLaVA v1.5, one possible reason is that ShareGPT4V uses a high-quality image-text dataset featuring highly descriptive captions for training, which makes it describe more entities shown in images. GPT-4o, while top-performing, still misses many entities, suggesting room for improvement even on recognition capability.
@@ 3.3.2 Cognition
Table 3 shows Cognition Scores of LVLMs on Description task. Similarly, GPT-4o achieves the best performance and there is a large performance gap between GPT-4 and other open-source models. For open-source models, Qwen-VL-Chat achieves the best performance. In terms of different reasoning capabilities, all of the LVLMs show better performance on [Location Reasoning] than others, which is probably because it is a kind of relatively lower-level reasoning. Differently, for [Event Reasoning], [Event Relationship Reasoning], and [Next Moment Event Reasoning], all of the open-source LVLMs show very low performance, indicating they almost do not understand the story in the images at all. The significantly large performance gap on these three kinds of reasoning types between GPT-4 and open-source LVLMs could be a manifestation of the capabilities emerging in GPT-4. Though GPT-4o achieves the best performance, there is also a huge gap between its ability and the Oracle, and the gap is obviously larger than that of recognition scores. This indicates that LVLMs still have a lot of room for development in terms of cognitive abilities.
Table 3: Cognition Scores of LVLMs on Description task evaluated by GPT-4. For reference, the Cognition Score of Oracle is calculated based on the annotated [Description] in CogBench as an estimated upper bound.
InstructBLIP-7B 0.15 0.54 0.25 0.30 0.10 0.05 0.02 0.17 0.17
Qwen-VL-Chat 0.28 0.57 0.28 0.26 0.14 0.12 0.04 0.13 0.19
LLaVA-v1.5-7B 0.09 0.45 0.15 0.18 0.09 0.05 0.02 0.13 0.13
LLaVA-v1.5-13B 0.13 0.49 0.18 0.24 0.10 0.05 0.04 0.16 0.15
mPLUG-Owl-2 0.06 0.48 0.24 0.21 0.08 0.04 0.02 0.14 0.14
ShareGPT4V-7B 0.19 0.60 0.21 0.22 0.10 0.04 0.03 0.15 0.16
ShareGPT4V-13B 0.23 0.57 0.24 0.26 0.12 0.07 0.03 0.14 0.17
GPT-4V 0.43 0.71 0.47 0.24 0.35 0.31 0.10 0.46 0.37
GPT-4o 0.51 0.77 0.55 0.37 0.44 0.39 0.15 0.40 0.43
Oracle 0.92 0.98 0.94 0.81 0.98 0.92 0.90 0.92 0.93
3.3.3 Case Study of Description Task
Figure 5 shows a case of Description task. In terms of recognition, GPT-4V shows a good performance by recognizing most annotated entities such as boy, girl, man, woman, table, dress, hat, suit and only fails to recognize sofa, television, handbag . LLaVA-v1.5-13B obviously recognizes fewer entities than GPT-4V, as it only recognizes boy, woman, dress. However, GPT-4V fails to understand the story in the image and gets a 0 in terms of cognition. This is because it neither clarifies the location nor the relationships between the characters, nor does it accurately depict the events occurring in the image, etc. LLaVA-v1.5-13B is slightly better and successfully inferred that the woman is the mother. This case demonstrates that CogBench reveals current LVLMs falling short in recognition and cognition, with a gap remaining between their cognitive abilities and human levels.
@@ 3.3.3 Case Study of Description Task
Figure 5: Case study of the Description task. A representative open-source LVLM LLaVA-v1.5-13B (red frame) and GPT-4V (green frame) are selected for analysis. Recognized entities are marked in blue, and CoRs are marked in green.
3.3.4 Effectiveness Analysis of GPT-based Cognition Evaluation of Description Task
Table 4: CoR accuracy of evaluation methods.
@@ 3.3.4 Effectiveness Analysis of GPT-based Cognition Evaluation of Description Task
To validate GPT-based cognition evaluation, we manually scored the CoRs of 20 images with a binary scale (0/1) and compared various evaluation methods’ accuracy on this subset. Table 4 reveals that GPT-4 offers the highest accuracy, demonstrating that GPT-based evaluation aligns well with human assessment. Therefore it is effective to assess LVLMs’ performance on the Description task. Implementation details of evaluation methods beyond ChatGPT and GPT-4 can be found in Appendix F.
3.4 Results of VQA Task
Table 5: Model performance on VQA task. Each QA contains four options, with a chance rate of 25%.
@@ 3.4 Results of VQA Task
InstructBLIP-7B 0.56 0.65 0.49 0.48 0.36 0.38 0.44 0.43 0.44
Qwen-VL-Chat 0.65 0.82 0.60 0.54 0.51 0.45 0.47 0.51 0.54
LLaVA-V1.5-7B 0.58 0.81 0.54 0.55 0.46 0.45 0.54 0.53 0.53
LLaVA-V1.5-13B 0.70 0.82 0.65 0.60 0.50 0.47 0.58 0.57 0.57
mPLUG-Owl-2 0.51 0.82 0.59 0.55 0.46 0.48 0.47 0.52 0.53
ShareGPT4V-7B 0.58 0.80 0.64 0.54 0.49 0.40 0.51 0.54 0.53
ShareGPT4V-13B 0.67 0.80 0.65 0.56 0.49 0.50 0.6 0.55 0.56
GPT-4V 0.66 0.84 0.73 0.62 0.63 0.63 0.70 0.69 0.67
GPT-4o 0.78 0.91 0.77 0.73 0.75 0.73 0.83 0.77 0.77
Human 0.99 0.96 0.99 0.94 0.96 0.96 0.96 0.93 0.95
Table 5 shows the performance of LVLMs on VQA task. Consistent with results in previous sections, GPT-4o achieves the best performance and there is a performance gap between GPT-4o and GPT-4V. There is also a performance gap between GPT-4 and open-source models. For open-source models, ShareGPT4V-13B and LLaVA-v1.5-13B achieves better performance than other LVLMs based on 7B LLMs, which shows the importance of LLM size to LVLMs on this task. As for LVLMs based on 7B LLMs, the performance of InstructBLIP is the worst and there is also a gap between InstructBLIP and other models. One possible reason is that it only finetunes the Q-former for instruction-tuning as Q-former has limited capacity compared with LLMs. The performance of other models are similar. Consistent with previous findings, reasoning about location is also the easiest for LVLMs and event-related reasoning are more difficult. There is also a large gap between the performance of LVLMs and human beings. Note that the accuracy of Human in Table 5 is calculated based on the responses of five healthy people. They all have obtained a bachelor’s degree and are between the ages of 20 and 30.
@@ 4 Related Work
To better understand emerging capabilities of LVLMs, many scholars have proposed different evaluation benchmarks. LVLM-eHub [24] evaluates LVLMs’ multimodal capabilities across six categories using various publicly available computer vision datasets. MME [10], MMBench [11] and SEED Bench [12] use True/False Questions or Multiple-Choice Questions to evaluate different abilities of LVLMs. MM-VET [25] evaluates LVLMs in terms of their integrated VL capabilities via open-ended questions. Different from them, CogBench mainly focus on high-level cognition evaluation of LVLMs. Though some of them also consider cognition as one of the evaluation dimensions, most images they use can only be used to evaluate limited aspects of reasoning.
@@
Image Captioning is a classical VL task. As model capabilities advance, researchers strive to enhance their ability to describe images in detail. Krause et al. [26] propose Image Paragraph Captioning, tasking models with generating a descriptive paragraph for an image. Recently, some researchers [17, 18, 19, 23] are trying to leveraging the ability of LLMs to generate more detailed image descriptions. None of these tasks consider to evaluate the high-level cognitive ability of models through description. HL dataset [27] requires models to generate high-level captions, but it only considers three aspects (scene, action, rationale). The content of most images in existing datasets does not reach the level of a story. This reveals the need for higher-level datasets like CogBench.
@@
Visual Reasoning is closely related to the cognitive ability of models. VCR [28] tasks models with answering visual questions using commonsense reasoning and justifying their answers. VisualCOMET [29] is a framework of visual commonsense reasoning tasks to predict past, future events, and present intents. Hessel et al. [30] utilize images from VCR and Visual Genome [31] to evaluate the ability of models to perform abductive reasoning. Fu et al. [32] propose a task to identify the time and location of a given image. CURE [33] is proposed to measure both the zero-shot reasoning performance and consistency of VLMs. Similarly, compared CogBench, these tasks considers less kinds of reasoning and CogBench can be seen as the next step of these efforts.
@@ 5 Conclusion
In this paper, we incorporated the idea of Cookie Theft picture description task into the evaluation of the high-level cognitive abilities of LVLMs and designed a novel evaluation benchmark called CogBench. Images in CogBench are of high quality and contain rich cognitive reasoning, which makes it different from existing image datasets. Experiments show that there is still a large gap between the cognitive ability of LVLMs and human beings, indicating CogBench is a challenging benchmark.

### v4

@@ 3.3.1 Recognition
Table 2 shows the Recognition Scores of models on the Description task. InternVL2-26B, LLaVA-OV-7B-Chat, GPT-4o and Qwen2-VL achieve relatively better performance than other models, which means they can recognize and describe more entities. GPT-4V, CogVLM and CogVLM2 also demonstrate competitive performance. It can be concluded that the recognition ability of open-source LVLMs is approaching that of GPT-4o, with some models even surpassing it. Nevertheless, other open-source LVLMs have a significant gap in recognition capability compared to GPT-4, indicating some LVLMs still have room for development before reaching the recognition capability of GPT-4. Note that some models perform worse in mode 2 than in mode 1, which could be because these models focus more on high-level semantic reasoning and overlook the description of some low-level entities. Although many LVLMs perform well, such as InternVL2-26B and GPT-4o, they still miss many entities, suggesting room for improvement in recognition capability.
@@
Mode 1 Mode 2
InstructBLIP-7B 40.0 36.4
Qwen-VL-Chat 43.3 45.8
LLaVA-v1.5-7B 39.8 41.2
LLaVA-v1.5-13B 41.0 39.3
mPLUG-Owl-2 37.4 37.8
ShareGPT4V-7B 46.9 47.3
ShareGPT4V-13B 48.7 47.4
LLaVA-v1.6-vicuna-7B 49.3 51.1
LLaVA-v1.6-vicuna-13B 53.3 53.9
LLaVA-v1.6-34B 52.2 52.0
CogVLM-Chat 61.6 56.5
CogVLM2-Llama3-Chat 62.3 58.9
InternVL2-26B 70.7 65.9
Qwen2-VL-7B 66.3 59.6
LLaVA-OV-7B-Chat 67.4 72.3
GPT-4V 62.9 56.5
GPT-4o 66.8 65.1
Oracle 92.8
Table 2: Recognition Scores of LVLMs on the Description task. Mode 1 and Mode 2 refer to the Spontaneous Description mode and the Directed Reasoning mode respectively. For reference, the Recognition Score of Oracle is calculated based on the annotated [Description] in CogBench dataset as an estimated upper bound. Numbers are presented in % with a full score of 100%.
InstructBLIP-7B 14.9 / 17.0 54.2 / 59.9 24.5 / 29.2 30.0 / 41.1 10.3 / 9.3 4.9 / 5.6 1.9 / 6.5 17.3 / 22.8 16.7 / 19.8
Qwen-VL-Chat 23.4 / 25.5 57.1 / 58.8 28.3 / 31.1 29.3 / 50.6 15.1 / 19.4 12.7 / 13.2 3.7 / 9.3 12.0 / 25.7 19.3 / 26.3
LLaVA-v1.5-7B 8.5 / 19.1 45.2 / 59.3 15.1 / 18.9 18.3 / 40.7 8.8 / 8.6 4.9 / 5.6 1.9 / 1.9 13.2 / 21.8 12.8 / 18.6
LLaVA-v1.5-13B 12.8 / 14.9 48.6 / 55.9 17.9 / 19.8 24.3 / 39.5 9.8 / 9.3 4.7 / 5.9 3.7 / 4.7 16.3 / 21.3 15.0 / 18.5
mPLUG-Owl-2 6.4 / 12.8 48.0 / 57.1 23.6 / 22.6 21.3 / 41.8 8.4 / 9.4 4.0 / 4.5 1.9 / 3.7 13.7 / 17.3 13.6 / 17.9
ShareGPT4V-7B 19.1 / 14.9 60.5 / 58.8 20.8 / 25.5 22.4 / 35.0 10.0 / 12.6 4.0 / 6.4 2.8 / 3.7 15.1 / 17.5 15.6 / 18.8
ShareGPT4V-13B 23.4 / 17.0 57.1 / 60.5 23.6 / 29.2 26.2 / 37.6 12.4 / 12.4 6.6 / 7.3 2.8 / 3.7 14.4 / 16.1 17.1 / 19.3
LLaVA-v1.6-vicuna-7B 17.0 / 25.5 61.0 / 62.1 23.6 / 29.2 25.1 / 47.9 12.1 / 14.1 7.5 / 8.5 2.8 / 8.4 17.0 / 23.0 17.7 / 23.1
LLaVA-v1.6-vicuna-13B 17.0 / 27.7 63.3 / 65.5 27.4 / 33.0 23.6 / 42.2 15.1 / 15.8 9.6 / 9.9 0.9 / 2.8 17.7 / 26.4 19.3 / 24.1
LLaVA-v1.6-34B 25.5 / 27.7 65.5 / 68.4 28.3 / 47.2 22.8 / 48.7 16.8 / 18.1 10.8 / 13.4 3.7 / 10.3 16.3 / 30.5 20.2 / 28.3
CogVLM-Chat 29.8 / 46.8 75.1 / 72.3 40.6 / 50.0 23.6 / 45.2 28.1 / 29.1 21.6 / 27.3 4.7 / 7.5 27.1 / 33.3 29.4 / 35.2
CogVLM2-Llama3-Chat 34.0 / 31.9 73.4 / 75.7 46.2 / 49.1 24.7 / 41.1 28.4 / 32.2 26.8 / 32.5 5.6 / 8.4 30.2 / 40.0 31.4 / 37.9
InternVL2-26B 40.4 / 44.7 77.4 / 79.7 48.1 / 54.7 27.8 / 54.0 27.4 / 31.5 24.5 / 27.8 3.7 / 15.0 36.7 / 44.1 32.7 / 40.2
Qwen2-VL-7B 38.3 / 34.0 72.3 / 75.1 48.1 / 48.1 19.4 / 53.2 25.1 / 23.8 20.0 / 20.2 4.7 / 9.3 29.5 / 33.3 28.4 / 33.1
LLaVA-OV-7B-Chat 42.6 / 46.8 75.1 / 76.3 50.9 / 48.1 18.3 / 46.4 25.5 / 30.4 25.6 / 26.4 9.3 / 17.8 33.8 / 42.4 30.9 / 37.9
GPT-4V 40.4 / 40.4 74.0 / 70.1 44.3 / 49.1 34.2/57.4 28.0 / 33.2 25.4 / 29.2 10.3 / 18.7 42.7 / 51.6 34.8 / 41.8
GPT-4o 38.3 /51.1 79.7/82.5 50.9/62.3 29.3/73.4 42.2/48.8 42.1/50.6 11.2/34.6 43.2/58.5 42.7/56.5
Oracle 91.5 97.8 94.3 81.4 98.1 92.2 89.7 92.3 93.2
Table 3: Cognition Scores of LVLMs on the Description task evaluated by GPT-4. The results of the evaluation under the Spontaneous Description mode and Directed Reasoning mode are presented before and after the “/” in each table cell. For reference, the Cognition Scores of Oracle are calculated based on the annotated [Description] in CogBench dataset as an estimated upper bound. Numbers are presented in % with a full score of 100%.
@@ 3.3.2 Cognition
General analysis. Table 3 shows the Cognition Scores of LVLMs on the Description task. GPT-4o achieves the best performance and there is a large performance gap between GPT-4o and other open-source models. Among open-source models, InternVL2-26B, LLaVA-OV-7B-Chat, CogVLM models and Qwen2-VL achieve relatively better performance, with some approaching that of GPT-4V. In terms of different reasoning capabilities, all LVLMs show better performance on [Location Reasoning] than others, probably because it is a kind of relatively lower-level reasoning. Differently, for [Event Reasoning], [Event Relationship Reasoning], and [Next Moment Event Reasoning], most open-source LVLMs show lower performance. The Cognition Scores of some open-source LVLMs are only around or even lower than 10% across these dimensions, indicating they almost do not understand the story in the images. In contrast, GPT-4o shows significantly better performance in these three kinds of reasoning capabilities. Besides, though GPT-4o achieves the best performance, there is also a large gap between its Cognition Scores and the Oracle scores. This indicates that LVLMs still have a lot of room for development in terms of cognitive abilities.
Analysis based on different evaluation modes. When the Directed Reasoning mode is applied, significant performance improvements in the LVLMs can be observed compared to the Spontaneous Description mode. This suggests that current LVLMs, while potentially grasping some high-level semantic information in images, cannot spontaneously generate comprehensive descriptions through reasoning from various perspectives without appropriate prompts. Another intriguing finding is that for some open-source LVLMs, shifting from the Spontaneous Description mode to the Directed Reasoning mode results in noticeable performance improvements across various dimensions, such as [Character Relationship Reasoning]; however, for [Event Reasoning] and [Event Relationship Reasoning], the enhancements are rather limited. This further highlights their struggle to fully comprehend events in the images, with event-based reasoning being a key area where open-source LVLMs differ significantly from GPT-4o.
Analysis of recognition and cognition abilities. The performance gap between GPT-4o and open-source LVLMs in terms of cognition is obviously larger than that of recognition. This suggests that cognitive reasoning is a more advanced ability than recognition, and that these LVLMs still have significant room for development.
3.3.3 Case Study
@@
Figure 5: Case study of the Description task. The description is generated by GPT-4o in the Directed Reasoning mode. Recognized entities are marked in blue, and CoRs are marked in green.
Figure 5 shows a failure case of GPT-4o on the Description task under the Directed Reasoning mode. In terms of recognition, GPT-4o shows a good performance by recognizing most annotated entities such as men, women, buildings, phone booth, newspaper, coat, snow, sign and only fails to recognize bus stop and wind. However, GPT-4o fails to understand the story in the image and gets a 30.0% in terms of cognition. One of the most important reasons is that it does not recognize that the man is in the phone booth to escape the cold, rather than to make a phone call. This case demonstrates that CogBench reveals current LVLMs falling short in cognition, with a gap remaining between their cognitive abilities and human levels.
3.3.4 Effectiveness of GPT-based Evaluation
To validate the GPT-based cognition evaluation method, we manually scored CoRs of 20 images with a binary scale (0/1) and compared the accuracy of various evaluation methods on this subset. Table 4 reveals that GPT-4 offers the highest accuracy, demonstrating that GPT-based evaluation aligns well with human assessment. Therefore, it is effective to assess LVLMs’ performance on the Description task. Implementation details of evaluation methods beyond ChatGPT and GPT-4 (Lin, 2004; Zhang et al., 2020; Sellam et al., 2020; He et al., 2021; Yin et al., 2021) can be found in Appendix F.
@@ 3.3.4 Effectiveness of GPT-based Evaluation
Table 4: CoR accuracy of cognition evaluation methods for the Description task.
3.4 Results of the VQA Task
@@
InstructBLIP-7B 60.5 71.4 48.8 54.9 40.3 36.8 46.7 47.6 47.4
Qwen-VL-Chat 65.1 82.3 60.5 54.3 50.9 45.0 47.4 51.1 54.0
LLaVA-V1.5-7B 58.1 81.4 54.3 54.6 45.9 45.0 54.1 52.6 52.8
LLaVA-V1.5-13B 69.8 82.3 65.4 59.9 50.2 47.3 57.8 57.1 57.3
mPLUG-Owl-2 51.2 81.8 58.6 54.6 46.0 47.5 47.4 51.8 52.7
ShareGPT4V-7B 58.1 80.5 63.6 53.6 48.8 40.3 51.1 54.1 53.4
ShareGPT4V-13B 67.4 80.0 65.4 56.5 49.4 49.8 60.0 54.6 56.3
LLaVA-v1.6-vicuna-7B 60.5 80.9 57.4 55.8 51.4 43.8 55.6 58.6 55.8
LLaVA-v1.6-vicuna-13B 66.3 85.9 64.8 59.3 58.2 53.7 65.9 60.0 61.5
LLaVA-v1.6-34B 80.2 92.7 83.3 74.8 68.4 66.9 68.8 74.2 73.7
CogVLM-Chat 73.3 87.7 75.9 66.9 58.5 53.2 63.0 63.7 64.2
CogVLM2-Llama3-Chat 73.3 92.3 86.4 76.0 71.9 62.9 67.4 71.9 73.5
InternVL2-26B 81.4 91.8 82.7 78.2 71.9 67.2 66.7 72.0 74.4
Qwen2-VL-7B 80.2 90.0 82.1 75.1 66.4 63.2 72.6 69.8 71.6
LLaVA-OV-7B-Chat 81.4 93.2 85.2 75.1 71.6 69.4 71.9 74.5 75.4
GPT-4V 70.9 81.8 72.8 63.7 63.4 66.9 68.9 69.2 68.0
GPT-4o 81.4 90.9 81.5 68.5 75.5 74.1 83.7 77.1 77.1
Human 98.8 95.9 98.8 94.3 95.6 96.0 96.3 93.3 95.3
Table 5: Model performance on the VQA task. Each QA contains four options, with a chance rate of 25%. Numbers are presented in % with a full score of 100%.
Table 5 shows the performance of LVLMs on the VQA task. GPT-4o achieves the best performance. Among open-source LVLMs, LLaVA-OV-7B-Chat, InternVL2-26B, LLaVA-v1.6-34B, CogVLM2, and Qwen2-VL-7B demonstrate better results, approaching the level of GPT-4o.
Consistent with previous findings, reasoning about location is also the easiest for LVLMs and event-related reasoning dimensions are more difficult. There is also a large gap between the performance of LVLMs and humans. Note that the accuracy of Human in Table 5 is calculated based on the responses of five healthy people. They all have obtained a bachelor’s degree and are between the ages of 20 and 30. Furthermore, some LVLMs perform differently on the two tasks, e.g. LLaVA-v1.6-34B, which highlights the necessity of including both tasks in the design of CogBench.
@@ 4 Related Work
To better understand the emerging capabilities of LVLMs, various evaluation benchmarks have been proposed. LVLM-eHub Xu et al. (2023) evaluates LVLMs’ capabilities across six categories using various publicly available CV datasets. MME Fu et al. (2023), MMBench Liu et al. (2023c) and SEED Bench Li et al. (2023a) use True/False questions or multiple-choice questions to evaluate different abilities. MM-VET Yu et al. (2024) evaluates LVLMs in terms of their integrated VL capabilities with open-ended questions. In contrast, CogBench focuses on high-level cognitive evaluation. Though some of them also consider cognition as one of the evaluation dimensions, they do not provide a comprehensive evaluation and most images they use evaluate only limited aspects of reasoning.
@@
Image Captioning is a classical VL task Zhou et al. (2022). As model capabilities advance, researchers strive to enhance their ability to describe images in detail. Krause et al. (2017) propose Image Paragraph Captioning, tasking models with generating a descriptive paragraph for an image. Recently, some researchers Xie et al. (2022); Zhu et al. (2023a); Zhuge et al. (2023); Chen et al. (2023a) have been leveraging the ability of LLMs to generate more detailed image descriptions. None of these tasks considers evaluating the high-level cognitive abilities of models through description. HL dataset Cafagna et al. (2023) requires models to generate high-level captions, but it only considers three aspects (scene, action, rationale). The content of most images in existing datasets does not reach the level of a story. This reveals the need for higher-level datasets like CogBench.
@@
Visual Reasoning task is closely related to the cognitive abilities of models. Visual Commonsense Reasoning (VCR) Zellers et al. (2019) tasks models with answering visual questions using commonsense reasoning and justifying their answers. VisualCOMET Park et al. (2020) is a framework of visual commonsense reasoning tasks to predict past, future events, and present intents. Hessel et al. (2022) utilize images from VCR and Visual Genome Krishna et al. (2017) to evaluate the ability of models to perform abductive reasoning. Fu et al. (2022) propose a task to identify the time and location of a given image. CURE Chen et al. (2023b) is proposed to measure both the zero-shot reasoning performance and consistency of VLMs. Similarly, these tasks consider fewer kinds of reasoning, and CogBench can be seen as the next step of these efforts.
@@ 5 Conclusion
In this paper, we incorporate the concept of the Cookie Theft picture description task into the evaluation of high-level cognitive abilities of LVLMs and design a novel evaluation benchmark called CogBench. The images in CogBench are of high quality and require more complex cognitive reasoning for interpretation, distinguishing it from existing image datasets. Experiments reveal a significant gap between the cognitive abilities of LVLMs and humans, highlighting CogBench as a challenging benchmark.
Limitations
Given the scarcity of images that meet our data collection standards, the number of images in CogBench is relatively limited. However, the number of images in CogBench is comparable to existing evaluation benchmarks, such as MM-VET, and is sufficient to serve as a reliable evaluation benchmark.
Ethical Considerations
Most images in CogBench are manually collected from Pinterest in accordance with its terms of service. The images are used under fair use for research purposes only and we will share our dataset with researchers who adhere to the ethical considerations established in this study. During the annotation process, we ensure that our annotators receive a fair wage and promptly address any questions they have. They are free to take breaks or quit the annotation task at any time.
Acknowledgments
This work has been supported by the China NSFC Project (No. U23B2018). Kenny Q. Zhu was partly supported by NSF Award No. 2349713.